On their way home from the amusement park, the gir
Answer #1

here,

radius , r = 65 m

us = 0.75

speed of car , v = 20.1 m/s

centripital force , fc = m* v^2 /r = 1250 * 20.1^2 /65 = 7769.4 m/s^2

the frictional force , ff = uk * m * g

ff = 0.75 * 1250 * 9.8 = 9187.5 N

as ff > fc

the car will stay on the road

for the safe speed

Fc = ff

m * u^2 /r = 9187.5

1250 * u^2 /65 = 9187.5

u = 21.9 m/s

so,

the correct option is B)

9)

mass , m = 8 kg

x = 4.28 cm = 0.0428 m

the force constant of spring , k = F/x

k = m * g /x = 9.8 * 8 /0.0428 N/m

k = 1832 N/m

the spring constant is A) 1832 N/m